### v3

@@
Impact of Leakage on Data Harmonization in Machine Learning Pipelines in Class Imbalance Across Sites.
Nicolás Nieto Institute of Neuroscience and Medicine (INM-7: Brain and Behaviour), Research Centre Jülich, Jülich, Germany Institute of Systems Neuroscience, Heinrich Heine University Düsseldorf, Düsseldorf, Germany Department of Cardiology, Pulmonology and Vascular Medicine, University Hospital and Medical Faculty, Heinrich-Heine University, Duesseldorf, Germany Correspondence: n.nieto@fz-juelich.de Simon B. Eickhoff Institute of Neuroscience and Medicine (INM-7: Brain and Behaviour), Research Centre Jülich, Jülich, Germany Institute of Systems Neuroscience, Heinrich Heine University Düsseldorf, Düsseldorf, Germany Christian Jung Department of Cardiology, Pulmonology and Vascular Medicine, University Hospital and Medical Faculty, Heinrich-Heine University, Duesseldorf, Germany Cardiovascular Research Institute Düsseldorf (CARID), Medical Faculty, Heinrich-Heine University, Duesseldorf, Germany Martin Reuter Artificial Intelligence in Medical Imaging, German Center for Neurodegenerative Diseases (DZNE), Bonn, Germany Department of Radiology, Harvard Medical School, Boston, MA, USA Kersten Diers Artificial Intelligence in Medical Imaging, German Center for Neurodegenerative Diseases (DZNE), Bonn, Germany for the Alzheimer’s Disease Neuroimaging Initiative Data used in preparation of this article were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database (adni.loni.usc.edu). As such, the investigators within the ADNI contributed to the design and implementation of ADNI and/or provided data but did not participate in analysis or writing of this report. A complete listing of ADNI investigators can be found at: (http://adni.loni.usc.edu/wp-content/uploads/how_to_apply/ADNI_Acknowledgement_List.pdf) Malte Kelm Department of Cardiology, Pulmonology and Vascular Medicine, University Hospital and Medical Faculty, Heinrich-Heine University, Duesseldorf, Germany Cardiovascular Research Institute Düsseldorf (CARID), Medical Faculty, Heinrich-Heine University, Duesseldorf, Germany Artur Lichtenberg Department of Cardiac Surgery, University Hospital and Medical Faculty, Heinrich-Heine University, Duesseldorf, Germany Federico Raimondo Institute of Neuroscience and Medicine (INM-7: Brain and Behaviour), Research Centre Jülich, Jülich, Germany Institute of Systems Neuroscience, Heinrich Heine University Düsseldorf, Düsseldorf, Germany Kaustubh R. Patil Institute of Neuroscience and Medicine (INM-7: Brain and Behaviour), Research Centre Jülich, Jülich, Germany Institute of Systems Neuroscience, Heinrich Heine University Düsseldorf, Düsseldorf, Germany
Summary
Machine learning (ML) models benefit from large datasets. Collecting data in biomedical domains is costly and challenging, hence, combining datasets has become a common practice. However, datasets obtained under different conditions could present undesired site-specific variability. Data harmonization methods aim to remove site-specific variance while retaining biologically relevant information. This study evaluates the effectiveness of popularly used ComBat-based methods for harmonizing data in scenarios where the class balance is not equal across sites. We find that these methods struggle with data leakage issues. To overcome this problem, we propose a novel approach “PrettYharmonize”, designed to harmonize data by pretending the target labels. We validate our approach using controlled datasets designed to benchmark the utility of harmonization. Finally, using real-world MRI and clinical data, we compare leakage-prone methods with “PrettYharmonize” and show that it achieves comparable performance while avoiding data leakage, particularly in site-target-dependence scenarios.
Keywords
Data Harmonization, ComBat, Data leakage, Machine learning, Medical Imaging, Magnetic Resonance Imaging, medical AI, clinical, ICU.
@@ Introduction
Many research fields have greatly benefited from machine learning (ML) approaches. ML models can extract important values from large amounts of data. Having vast data benefits the model’s classification performance and helps capture the underlying patterns, promoting better generalization to new unseen data. This makes combining multiple datasets an appealing approach, especially in domains where obtaining data in a uniform setting is challenging 1. Moreover, small health or research centers that can not afford to collect a large number in-house data, using data acquired in different sites is the only possibility for train ML models. However, different datasets obtained under different conditions often present variability due to differences in the acquisition procedure that are unrelated to relevant biological information 2. This undesired variability, also known as Effects of Site (EoS), can induce biased results if present or not correctly removed 3. These differences may come from systematic differences, which can be corrected, or random variations, which can not be modeled or corrected by harmonization. This problem is of common occurrence in many biomedical domains. For example, clinical data is affected by the acquisition site, as different hospitals have different laboratory machines, procedures, and criteria. Another example is the medical imaging field, as images are affected by acquisition protocol, scanner drifts, and time of the day, just to name a few factors 3, 4. Within this field, Magnetic Resonance Imaging (MRI) images are particularly susceptible to this site-related variance, like the magnetic field strength, room temperature fluctuation or changes in the electromagnetic noise, which makes even images obtained from scanners with the same manufacturer and the same parameters exhibit different characteristics 5, 6. Many works showed that removing this undesired systematic variability, which is only related to the acquisition site and has no biological information, can benefit further analysis made with the data 7, 8, 9, 10, 11. To this end, several Methods Aiming to Remove the Effects of Site (MAREoS) have been proposed and developed 4, 12. These MAREoS methods are typically used as a pre-processing step, where the site effects are removed and the “site-effect free” data, also known as harmonized data, is used for statistical analysis or to train and evaluate ML models.
Among these MAREoS, the ones based on “ComBat” are extensively used in several domains. The ComBat method was originally proposed for correcting batch differences in genomic data 13 and was later adapted to other domains like MRI data 14, 7. ComBat uses Bayesian regression to find additive (location) and multiplicative (scale) corrections for each feature in each site. Within the ComBat-based methods, “neuroHarmonize” was proposed 15 to allow for the preservation of non-linear covariate effects and has been widely used since 4, 12, 16, 17. Although ComBat and its derivations have been widely applied in medical imaging data, several concerns have been raised, mainly because ComBat’s hypothesis and assumptions only hold for genomic data, where it was originally proposed, and may not be fulfilled in other applications fields 18. Additionally, concerns had been raised on the integration of ComBat into ML models, as the location and scale parameters of the model could not be learned in a subset of data (train data) and then applied to a new unseen subset of data (test data) 19.
Early implementations of ComBat 14, 20 used the whole dataset to estimate the model’s parameters and create a harmonized dataset that is then used from all the downstream analyses. This approach was used in several works 7, 8, 21, 22, 23, 24. While this approach is valid when performing statistical analyses, it is not consistent with machine learning applications where the training and test data must be separate 25, 19. Specifically, the parameters of the models, including preprocessing models, must be obtained on a training set and then applied to the test set. This separation is important to get realistic estimates of generalization performance (e.g. using cross-validation) and to ensure deployability of the model in the real world where the test data is not yet available 26, 27.
ComBat-MAGA 28, neuroHarmonize 15, and ”harmonizer” 19, which is based on neuroHarmonize, allow the estimation of the model’s parameters in a training set and apply them to the test samples, however, a critical assumption of ComBat is that all variance not shared across sites is unwanted site-related variance. Consequently, ComBat removes any variance that is not common to all sites, including the relevant biological variance. This poses a new problem, as this assumption is broken when a class imbalance occurs across sites and a target-site dependence exists, for example when the control patients are acquired in one site and target patients in a different one 29. This could also be extended to other possible biological information like comorbidities or disease severity, for example, if more severe patients are consistently treated or acquired only in one site. In these cases, even though ComBat will provide harmonized data, it will remove the variance related to the target (control versus patient) as the assumption is that only non-relevant factors change between sites.
ComBat allows to retain the biologically relevant variables, for example, a diagnosis, the age, or the sex of a patient, by providing these variables as covariates to be retained. Nonetheless, this information is needed both when training the model and when applying the model to the test data. Thus, if target labels need to be preserved, this inevitably leads to the model requiring the test labels preventing the model’s use in real-world applications where test labels are not available or known18. This phenomenon where information of the test set is presented to the models is commonly known as data leakage. It is also well described that leaking the test target information would produce overconfident results, which could be misleading and can jeopardize the progress of an entire research field, as researchers who avoid data leakage would not be able to outperform the models that present data leakage 26, 27, 25.
@@
In this work, we aim to empirically demonstrate a shortcoming of ComBat-based harmonization in site-target dependence scenarios, i.e., that the model can properly harmonize the data only when test labels are used and data leakage happens. To do so, we performed controlled experiments for age regression and sex classification using real MRI data for healthy control individuals. Also using MRI data, a dementia and mild cognitive impairment (MCI) classification experiment was performed. Additionally, an outcome prediction of septic patients was performed using clinical data. All experiments were conducted both in site-target dependence and independence scenarios. Several harmonization schemes were used and compared, allowing and not allowing leakage, to harmonize the data.
Finally, to overcome the aforementioned problem, we propose a new harmonization method, called PRETended Target Y Harmonize (PrettYHarmonize), which allows the users to integrate ComBat in an ML pipeline, harmonizing the data and generating a prediction without using the test labels and thus avoiding data leakage. We validated our method using benchmark datasets 3. Additionally, the proposed method was compared with the other harmonization schemes on the site-target dependence and independence scenarios to comprehensively compare no harmonization, leakage, and no-leakage methods. The corresponding Python package is publicly available via GitHub https://github.com/juaml/PrettYharmonize.
@@ 0.1 PrettYharmonize validation
The proposed PrettYharmonize method is based on a neuroHarmonize model15 but uses the combination of two ML models, a Predictive and a Stack model to harmonize the data without using test labels, preventing data leakage by design. To avoid using the test data labels, the proposed methods rest on the use of pretended target values, which are used to harmonize the test data and generate a prediction with the Predictive model. Our main assumption is that when harmonizing the test data with the correct label, e.g. when the pretended label matches the real test label, the neuroHarmonize model will preserve the relevant information. On the contrary, when the pretended label doesn’t correspond with the real label, the all information will be removed, both effects of site and biological information, as it was harmonized under the wrong test label assumption. Then, the Predictive model will generates a more accurate and Using these predictions as input features, the Stack model generates a final unique prediction for each sample. As test labels are pretended and not used, predictions can be generated without requiring test target values. A detailed description of the method workflow is presented in the Method-PrettYharmonize section.
@@
PrettYharmonize was validated using the datasets specially designed to validate MAREoS 3. This dataset consists of eight internal datasets simulating eighteen MRI features (cortical thickness, cortical surface area, or subcortical volumes). Four datasets contain a “True” signal and four only contain an Effect of Site (“EoS”) signal related to a binary target, which the MAREoS should remove to avoid fraudulent classification performance. For each kind of signal, two ”Simple” and ”Interaction” datasets are proposed, for linear and no-linear relationships between the features and the target. For each internal dataset, 1000 samples coming from eight different sites are simulated. An extended data description of this dataset is available in the “Data-MAREoS” section. On the datasets that presented True signal and no EoS, a Baseline model (Random Forest), trained on the unharmonized data, obtained a balanced Accuracy (bACC) of around 80%, as expected (colum. This model also obtained a close to 80% bACC on the datasets that only contained the EoS signal, but this time fraudulently used the EoS signal to perform the classification (Table 1).
@@
The model successfully removed the EoS in all datasets which only present an EoS signal. Furthermore, in those datasets where only the True signal was presented, the method did not degrade the real signal while aiming to remove EoS, which in the True datasets are not presented (Table 1). Finally, we repeated this analysis to check robustness using three different Predictive models; Gaussian Process Classifier (GP), Support Vector Machine with Radial basis kernel (SVM), and least absolute shrinkage and selection operator (LASSO). These yielded similar results (Tables Supp 1, 2, and 3).
@@
0.2 Forced site-target dependence and independence.
@@
To investigate the impact of the harmonization when site and target are dependent or independent, different scenarios were generated by thoroughly sampling data from different datasets. To force site-target dependence, for each site, we retained the majority of samples from one class and a small number of samples from the other classes. For example, let’s assume a binary classification problem where only two sites are presented. In this case, from site A, mainly samples from class one were retained, and only a few samples from class two. For site B, an opposite sampling strategy was used, retaining mainly samples from class two and just a few from class one. The small number of samples from the minority class were retained to avoid singular matrices, which the algorithms cannot support. In this case, we hypothesize that the traditional harmonization will remove important information, as the biological variance is related (or dependent) to the sites, unless test labels are leaked to the harmonization model. In contrast, site-target independence scenarios were generated by retaining the same amount of samples for each class sampled from each site. We hypothesize that in this case, the harmonization scheme will not remove important information, as the biological variance is common to all sites.
@@
A total of seven datasets were used in our experiments. Five datasets [AOMID-ID1000, eNKI, CamCAN, SALD, and 1000Brains] containing MRI data from healthy control participants were sampled for age regression and sex classification problems. Using the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset 30, where the data was collected in different sites, a classification of healthy controls versus mild cognitive impairment (MCI) / dementia was performed. Finally, using a publicly available multi-site clinical dataset [eICU] 31, 32 was used to perform a a classification of hospital discharge status of septic patients, as this is an important and extensively studied problem in the literature 33, 34, 35. An extensive and detailed explanation of all the datasets and the sampling method used to force site-target dependence and independence is presented in Section Data description.
Five harmonization schemes were evaluated. The proposed PrettYharmonize model performs leakage-free harmonization while pretending labels (see section Experimental procedures-PrettYharmonize). Further, two schemes that induce data leakage but have been used in the literature were implemented. A “Whole Data Harmonization” (WDH) scheme, where the pooled data from all sites was used to train a neuroHarmonize model (based on ComBat) and to obtain a harmonized dataset before splitting the data into train and test folds. Additionally, a “Test Target Leakage” (TTL) scheme was used, where a neuroHarmonize model learned its parameters only on the training data while retaining target variance and thus requires the test labels to perform harmonization. Additionally, a scheme called “No Target” was proposed, where the harmonization model learns its parameters on the train data but without explicitly retaining the target variance, and thus the test labels were not used for test set harmonization. Finally, a baseline model without harmonization was implemented, where the pooled data were used unharmonized.

### v4

@@
Abstract
Due to the cost and complexity of data collection in biomedical domains, it is common practice to combine data sets from multiple sites to obtain large datasets required for machine learning. However, undesired site-specific variability presents challenges. Data harmonization aims to address this issue by removing site-specific variance while preserving biologically relevant information. We show that the widely used ComBat-based harmonization improvements are driven by data leakage due to illicit use of target information when class labels are imbalanced across sites, a common scenario in biomedical domains. We propose a novel approach, PrettYharmonize, which leverages subtle differences in data harmonized using different pretended target values. Using controlled benchmark datasets and real-world magnetic resonance imaging and clinical data, we demonstrate that our leakage-free PrettYharmonize method achieves performance comparable to leakage-prone methods. As such, PrettYharmonize is a viable method to combine data sets for machine learning applications.
keywords:
Data Harmonization , ComBat , Data Leakage , Machine learning , Medical Imaging , Magnetic Resonance Imaging , medical AI , clinical , ICU
journal: Neurocomputing
[…]
@@ Introduction
Many research fields have greatly benefited from machine learning (ML) approaches. ML relies on large datasets to learn generalizable models, as these datasets help capture robust underlying patterns. This makes combining multiple datasets particularly attractive, especially in domains where obtaining data from a single location is challenging. However, combining datasets remains a significant challenge, as datasets obtained from different locations, even with similar acquisition parameters, often contain variability unrelated to relevant biological information [9, 2, 3].
This undesired site-related variability may stem from systematic differences, which can be corrected, or random variations, which cannot be modeled or corrected. The systematic variability, known as Effects of Site (EoS), is prevalent in many biomedical domains and can lead to biased results if not properly addressed [10]. For example, clinical data are influenced by the acquisition site, as different hospitals use varying laboratory equipment, procedures, and criteria. Similarly, medical imaging data are affected by factors such as acquisition protocols, scanner drifts, and even the time of day [10, 1]. Magnetic Resonance Imaging (MRI) is particularly susceptible to EoS, as variability can arise from differences in magnetic field strength, room temperature fluctuations, or electromagnetic noise—even when using scanners from the same manufacturer with identical parameters [11, 12]. To address this issue, numerous Methods Aiming to Remove the Effects of Site (MAREoS) have been proposed [1, 13, 7]. MAREoS are typically employed as a preprocessing step to generate “site-effect-free” data while preserving biological information, resulting in harmonized data that can enhance subsequent statistical and ML analyses [16, 23, 51, 52, 5].
ComBat-based MAREoS are extensively used in several domains. Originally proposed to correct batch differences in genomic data [14], ComBat was later adapted to other domains, including MRI data [15, 16]. The method employs Bayesian regression to estimate additive (location) and multiplicative (scale) corrections for each feature across sites. In addition to removing site effects, ComBat can preserve the variance of biologically relevant variables when provided as covariates. The ComBat-GAM extension was developed to account for nonlinear covariate effects, and its associated “neuroHarmonize” software has been widely adopted [17]. Although concerns have been raised that the assumptions of ComBat, originally designed for genomic data, may not always hold for other data types [22], numerous empirical studies have demonstrated its effectiveness [1, 13, 20, 21].
Several studies using ComBat within ML pipelines have failed to separate training and test data, incorrectly harmonizing the entire dataset before downstream analysis [15, 16, 23, 25, 26, 59, 27, 58]. While this approach is valid for statistical analysis, it is inconsistent with ML principles, where data leakage can lead to invalid models and misleading interpretations if training and test data are not properly separated [28, 55, 54]. Failing to maintain this separation during harmonization can result in overly optimistic generalization performance estimates, particularly in cross-validation settings [50]. To address this, techniques have been developed to integrate ComBat into ML pipelines as a preprocessing step, where its parameters are learned on the training data and then applied to unseen test data [17, 19, 50].
Although these methods correctly split train and test folds, further challenges arise due to the violation of a critical ComBat assumption - all variance not shared across sites is considered unwanted site-related variance. This becomes particularly problematic when biologically relevant variance is associated with site differences, such as in a diagnostic classification task where one site predominantly includes patients and another healthy controls [29], leading to class imbalance across sites. ComBat may eliminate variance related to the target in such cases of site-target dependence, resulting in a harmonized dataset that yields fewer or null findings in subsequent analyses. To mitigate this issue, the target variable is often included as a covariate in ComBat, ensuring its variance is preserved during harmonization. However, while preserving relevant variance is not inherently problematic, specifying the target as a covariate introduces another form of data leakage, as the target values of the test set are required to apply the harmonization model but are unavailable in real-world scenarios [28]. This critical issue has not been thoroughly investigated until recently [50].
@@
In this work, we aimed to empirically demonstrate a limitation of ComBat-based harmonization in site-target dependence scenarios. To this end, we conducted controlled experiments for age regression and sex classification using real MRI data from healthy individuals. Additionally, we performed two clinically relevant tasks: dementia and mild cognitive impairment (MCI) detection using MRI data, and outcome prediction for septic patients using intensive care unit (ICU) data. To systematically evaluate the impact of site-target association, all experiments were conducted under both site-target dependence and independence conditions. Specifically, we investigated the performance of different harmonization schemes, both with and without preserving target variance, under these scenarios. Preserving target variance leads to data leakage; however, this approach is still commonly used in the literature.
Finally, to address the data leakage issue in ML pipelines while combining data from multiple sites, we propose a novel method called “PRETended Target Y Harmonize” (PrettYHarmonize). Using a stacking ensemble model, PrettYHarmonize learns subtle differences in data harmonized with varying target values. It avoids data leakage by employing pretended target values for the test data. We validated our method using benchmark datasets [10] and demonstrate that PrettYHarmonize performs competitively compared to other harmonization schemes in both site-target dependence and independence scenarios. We provide a comprehensive comparison of no-harmonization, leakage, and no-leakage methods. The corresponding Python package of the proposed method is publicly available at https://github.com/juaml/PrettYharmonize, and the code to replicate our results is also publicly available at https://github.com/juaml/harmonize_project.
@@ 0.1 PrettYharmonize validation
The proposed PrettYharmonize method builds on the neuroHarmonize model [17] but introduces a novel approach by combining two machine learning models: a Predictive model and a Stack model. To harmonize test data without access to true labels, the method employs pretended test target values. These pretended values are used to harmonize the test data and generate predictions via the Predictive model. This ensures harmonization without using test labels, thereby preventing data leakage by design.
@@
Our key assumption is that when the test data is harmonized with the correct label (i.e., when the pretended label matches the true test label), the neuroHarmonize model will preserve biologically relevant information. Conversely, when the pretended label does not match the true label, both site-related effects and biological signals will be removed, as the harmonization is performed under an incorrect label assumption. The Predictive model generates predictions for each harmonized dataset, one for each class in the dataset, and the Stack model combines these predictions to produce a final, unified prediction for each sample. Since test labels are only pretended and not used, predictions can be generated without requiring access to the true test target values. Importantly, while incorporating a novel harmonization scheme internally, PrettYharmonize only uses the harmonized data internally and does not create a harmonized dataset as outcome, but rather a final target prediction. A detailed description of the method’s workflow is provided in the Method-PrettYharmonize section.
@@
To evaluate the harmonization capabilities of the harmonization scheme using pretended target of PrettYharmonize, a datasets specifically designed to benchmark MAREoS was used [10]. These datasets simulate eighteen MRI features, including cortical thickness, cortical surface area, and subcortical volumes, across eight internal datasets. Among these, four datasets contain a “True” signal, while the remaining four contain only an “EoS” signal related to a binary target. The EoS signal is designed to test whether MAREoS can remove it, thereby preventing fraudulent classification performance. For each type of signal (True and EoS), two variations are provided: “Simple” and “Interaction” datasets, representing linear and non-linear relationships between the features and the target, respectively. Each internal dataset comprises 1000 samples simulated from eight different sites. A detailed description of these datasets is available in the Data-MAREoS section.
On datasets containing the True signal and no EoS, a Baseline model (Random Forest) trained on unharmonized data achieved a balanced accuracy (bACC) of approximately 80%, as expected. However, the same model also achieved a bACC close to 80% on datasets containing only the EoS signal, fraudulently leveraging the EoS signal for classification (Table 1).
@@
Using Random Forest as both the Predictive and Stack models, PrettYharmonize successfully removed the EoS signal in all datasets containing only EoS. Furthermore, in datasets where only the True signal was present, the method preserved the real signal while aiming to remove EoS (Table 1). To ensure robustness, we repeated this analysis using three additional Predictive models: Gaussian Process Classifier (GP), Support Vector Machine with a Radial basis kernel (SVM), and Least Absolute Shrinkage and Selection Operator (LASSO). These models yielded similar results (Tables Supp 1, 2, and 3).
@@
0.2 Forced site-target dependence and independence scenarios.
To investigate the impact of harmonization in scenarios where the site and target are either dependent or independent, we systematically generated different experimental conditions by sampling data from various datasets.
@@
For site-target dependence, scenarios were created such that, for each site, the majority of samples were retained from one class, while only a small number of samples were retained from the other class(es). For example, in a binary classification problem with two sites, site A predominantly retained samples from class one, with only a few samples from class two, while site B mainly retained samples from class two, with a small number from class one. The inclusion of minority class samples was necessary to avoid singular matrices, which would otherwise cause computational issues for the algorithms. We hypothesize that traditional harmonization methods will remove important biological variance in this scenario unless test labels are leaked to the harmonization model, as the biological variance is intrinsically tied to the site distribution.
In contrast, for site-target independence, scenarios were generated by retaining an equal proportion of samples from each class across all sites. In this case, it is hypothesized that harmonization will not remove biologically relevant information, even when the target is not explicitly preserved as a covariate, as the biological variance is shared across all sites.
@@
A total of seven datasets were utilized in the experiments. Five MRI datasets (AOMIC-ID1000, eNKI, CamCAN, SALD, and 1000Brains), containing data from healthy control participants, were sampled for age regression and sex classification tasks. Additionally, the Alzheimer’s Disease Neuroimaging Initiative (ADNI) dataset [41], which includes data collected from multiple sites, was used to detect mild cognitive impairment (MCI) or dementia patients. Finally, the publicly available eICU clinical dataset [48, 47] was employed to classify the hospital discharge status of septic patients, a well-studied and clinically significant problem in the literature [42, 43, 44]. A comprehensive and detailed description of all datasets, as well as the sampling methods used to enforce site-target dependence and independence, is provided in the Data Description section.
Five harmonization schemes were evaluated in this study. Whole Data Harmonization (WDH) involves training a neuroHarmonize model on the pooled data from all sites to generate a harmonized dataset before splitting the data into training and test folds. This approach inherently leads to data leakage, as the test data is included in the harmonization training process [55, 50]. Another scheme, Test Target Leakage (TTL), trains the neuroHarmonize model on the training data while explicitly retaining target variance. Although this scheme does not use the test target to train the harmonization model, it requires test labels to transform (harmonize) the test data, resulting in data leakage. A third scheme, No Target, trains the harmonization model on the training data without explicitly retaining target variance. As a result, test labels are not used for harmonizing the test set, avoiding leakage but potentially removing biologically relevant information. A Unharmonized approach was implemented as a baseline, where the original data was pooled and used without any harmonization. The proposed PrettYharmonize model, which internally performs leakage-free harmonization, was benchmaked. By comparing these schemes, we aimed to evaluate the trade-offs between preserving biological signals, avoiding data leakage, and maintaining model performance.